Please help. I struggle with these.
In 2000, the circulation of a local newspaper decreased from 3041 papers sold in January to 2326 papers sold in December. Find the percent of decrease.

Answer by checkley77(12569)

(3041-2326)/3041=715/3041=.235=23.5% decrease.